OpenStudy (anonymous):

a long distance trucker traveled 72 miles in one direction during a snow storm. The return trip was accomplished in the rain at double speed in 2 hours less time. Find the speed going

OpenStudy (dumbcow):

d = r*t 72 = rt 72 = (2r)(t-2) set the equations equal rt = 2rt - 4r rt = 4r t =4 substitute 4 for t 72 = 4r 18 = r